A train left frog city traveling north at 60 mph. Two hours later, another train left frog city traveling south at 72 mph. How many hours had the first train been traveling when they were 450 miles apart?
Answer by rfer(12657)

a has 120 mi lead
450-120=330
330/132=2.5
2.5+2=4.5 hrs